OpenStudy (anonymous):

john drove 450 ml. in 8 hrs. hr drove 50 mph on the highway and 60mph on the turnpike. how many hrs did he drive on the turnpike?

OpenStudy (anonymous):

50x + 60(8-x) = 450 50x + 480 - 60x = 450 -10x +480-480 = 450-480 -10x /-10 = -30/-10 x=3 3 hours on the highway, 8-3=5 hours on the turnpike